# Controlling Space and Time with Diffusion Models

### v1

@@ 5 Experiments
We consider both in-distribution and OOD evaluations for 3D NVS capabilities in our ablations and comparisons to prior work. We use the RealEstate10K dataset (Zhou et al., 2018) as a common in-distribution evaluation. To maximize the amount of training data, we use 1% of the dataset as our validation split. We run inference on all baselines ourselves for this split, noting that they might instead be advantaged as our test data may exist in their training data (for PNVS, all our test data is in fact part of their training dataset, giving them a distinct advantage). For OOD evaluation, we use the LLFF dataset (Mildenhall et al., 2019).
| cRE10k test | |||||||||
| PNVS | 51.41 | 1007. | 248.9 | 0.9961 (1.000) | 0.9773 (1.024) | 0.3068 (0.3638) | 0.3899 | 16.07 | 0.3878 |
| MotionCtrl | 43.07 | 370.6 | 411.7 | 0.4193 (1.000) | 1.027 (1.032) | 0.2549 (0.3634) | 0.5003 | 12.74 | 0.2668 |
| 4DiM-R | 31.23 | 306.3 | 195.1 | 0.9974 (1.000) | 1.023 (1.075) | 0.3029 (0.3413) | 0.2630 | 18.09 | 0.5309 |
| 4DiM | 31.96 | 314.9 | 221.9 | 0.9935 (1.000) | 1.008 (1.034) | 0.3326 (0.3488) | 0.3016 | 17.08 | 0.4628 |
| cLLFF zero-shot | |||||||||
| PNVS | 185.4 | 2197. | 1751. | 0.7235 (0.9962) | 0.9346 (0.8853) | 0.2213 (0.1857) | 0.5969 | 12.04 | 0.1311 |
| MotionCtrl | 106.0 | 531.5 | 1754. | 0.1515 (0.9962) | 0.9148 (0.9415) | 0.1366 (0.2216) | 0.7016 | 9.722 | 0.06756 |
| 4DiM-R | 63.48 | 353.2 | 841.6 | 0.9659 (0.9962) | 0.9265 (0.8951) | 0.2011 (0.1934) | 0.5403 | 11.55 | 0.1444 |
| 4DiM | 63.78 | 356.8 | 864.4 | 0.9167 (0.9962) | 0.9131 (0.8960) | 0.1838 (0.1943) | 0.5415 | 11.58 | 0.1408 |
| RE10k test | |||||||||
| 4DiM-R (RE10k) | 32.40 | 335.9 | 212.0 | 1.000 (1.000) | 1.021 (1.022) | 0.3508 (0.3737) | 0.3002 | 16.64 | 0.4704 |
| 4DiM-R (cRE10k) | 31.23 | 306.3 | 195.1 | 0.9974 (1.000) | 1.023 (1.075) | 0.3029 (0.3413) | 0.2630 | 18.09 | 0.5309 |
| LLFF zero-shot | |||||||||
| 4DiM-R (RE10k) | 71.07 | 521.2 | 740.1 | 0.7727 (0.9962) | 0.9003 (0.8876) | 0.2108 (0.1915) | 0.4568 | 12.87 | 0.2016 |
| 4DiM-R (cRE10k) | 63.48 | 353.2 | 841.6 | 0.9659 (0.9962) | 0.9265 (0.8951) | 0.2011 (0.1934) | 0.5403 | 11.55 | 0.1444 |
| ScanNet++ zero-shot | |||||||||
| 4DiM-R (RE10k) | 23.68 | 248.0 | 137.9 | 0.9512 (0.9815) | 1.885 (1.858) | 1.534 (1.520) | 0.1938 | 20.74 | 0.6716 |
| 4DiM-R (cRE10k) | 22.89 | 243.2 | 130.6 | 0.9685 (0.9815) | 1.851 (1.917) | 1.541 (1.550) | 0.1809 | 21.25 | 0.6952 |
| cRE10k test | |||||||||
| 4DiM (no video) | 32.45 | 326.5 | 207.7 | 0.9974 (1.000) | 1.023 (1.029) | 0.3356 (0.3692) | 0.2713 | 17.80 | 0.5209 |
| 4DiM | 31.96 | 314.9 | 221.9 | 0.9935 (1.000) | 1.008 (1.034) | 0.3326 (0.3488) | 0.3016 | 17.08 | 0.4628 |
| ScanNet++ test | |||||||||
| 4DiM (no video) | 24.61 | 250.7 | 129.2 | 0.9615 (0.9815) | 1.895 (1.764) | 1.446 (1.431) | 0.1824 | 21.34 | 0.6969 |
| 4DiM | 23.48 | 223.7 | 146.0 | n/a (0.9815) | 1.706 (1.889) | 1.550 (1.503) | 0.1965 | 20.67 | 0.6434 |
| cLLFF zero-shot | |||||||||
| 4DiM (no video) | 64.34 | 401.4 | 884.1 | 0.9886 (0.9962) | 0.9157 (0.9067) | 0.1872 (0.1965) | 0.5437 | 11.45 | 0.1430 |
| 4DiM | 63.78 | 356.8 | 864.4 | 0.9167 (0.9962) | 0.9131 (0.8960) | 0.1838 (0.1943) | 0.5415 | 11.58 | 0.1408 |
@@
We present our main results on 3D novel view synthesis conditioned on a single image in Tab. 1 and Fig. 3, comparing the capabilities of our 4DiM cascade against PhotoConsistentNVS222For PNVS, we follow Yu et al. (2023a) and use a Markovian sliding window for sampling, as they find it is the stronger than stochastic conditioning (Watson et al., 2022). (Yu et al., 2023a) (PNVS) and MotionCtrl (Wang et al., 2023b). These two are the strongest 3D scene NVS diffusion models with code and checkpoints available. For LLFF, we load the input trajectory of views in order, subsample frames evenly, and then generate 7 views given a single input image. RealEstate10K trajectories are much longer, so following PNVS, we subsample with a stride of 10. We use the 7-view NVS task conditioned on a single image, in order to avoid weakening the baselines: MotionCtrl can only predict up to 14 frames, and PNVS performance degrades with the length of the sequence as it is an image-to-image model. We trained versions of 4DiM that process 8-frames to this end (as opposed to using the main 32-frame 4DiM model and subsampling the output) for a more apples-to-apples comparison. Quantitative metrics are computed on 128 scenes from our RealEstate10K test split, and on all 44 scenes in LLFF.
@@
We find that 4DiM achieves superior results in fidelity (FID, FDD, FVD) and metrc-scale pose alignment (LPIPS, PSNR, SSIM), by a wide margin compared to the baseline diffusion models. Qualitatively, we find that MotionCtrl has trouble aligning with the conditioning poses. We also observe that PNVS exhibits more artifacts, with a substantial drop in quality in the out-of-distribution setting. Surprisingly, PNVS achieves the best TSED in LLFF but upon closer inspection, we find that the number of keypoint matches is 3x less than for 4DiM. This suggests potential improvements for TSED, such as a higher threshold for the number of minimum keypoint matches, or requiring more spatial coverage of keypoints to discard examples for which the matches are too localized.
@@
One of the most critical ingredients for 3D NVS, especially in the out-of-distribution setting, is the training dataset. We ablate this in three ways:
@@
Table 1 shows that adding more diverse 3D data beyond RealEstate10K to the training mixture improves pose alignment (SfM distances) on LLFF (zero-shot). The evaluation suggests some loss in fidelity when using the full mixture which is expected since the mixture is more diverse and includes both indoor and outdoor scenes.
@@
To determine the impact of consistent, metric scale training data, we compare a model trained with on the original RE10k data against one with calibrated metric scale poses, with all else unchanged. Neither model includes the other 3D data sources in their training data used to train our best 4DiM model, because these are scale-calibrated and would confound this study otherwise. Quantitative results on 3D NVS from a single conditioning image are given in Tab. 2 (including zero-shot performance on LLFF and ScanNet++), and qualitative results are given in Fig. 4. We find that models trained on uncalibrated data often overshoot or undershoot when scale is ambiguous, and that models trained on scale-calibrated data solve this problem.
@@
We also show the importance of leveraging video as a training data source. Table 3 presents results comparing 8-frame 4DiM trained on our final data mixture to the otherwise identical model but trained without video data. We find that including video data improves fidelity and generalization ability which is expected since available 3D/4D data is not as diverse. The improvement is more apparent qualitatively (see Supplementary Material F for more samples).
@@ 5.2 Emergence of temporal dynamics
Early in our experiments, we observed that video from a single image can often yield samples with little motion. While this can be remedied by lowering the guidance weight on the conditioning image, this can entail a cost in fidelity. Nevertheless, with two or more input frames, 4DiM can successfully interpolate and extrapolate video. We quantitatively illustrate these results through our keypoint distance metric (along with other standard metrics) in Tab. 4. We suspect that future work including additional controls (e.g., text) on models like 4DiM, as well as larger scale, will likely play a key role in breaking modes for temporal dynamics, given that text+image-to-video has been successful for large models (Brooks et al., 2024).
| Input | ||||||
StreetView |
| Stitched using gamma adjustment and homography warp | ||||||
| Generated | ||||||
@@ 5.3 Multiframe Conditioning
While we (and most baselines) focus on single frame conditioning, using more than one frame unlocks interesting capabilities. Here we show two applications. First, we consider 4DiM on the task of panorama stitching (rendering novel viewpoints), given discrete images that collectively cover a full 360 FOV (see Fig. 5). For reference we stitch the images using a naive homography warp and gamma adjustment (Brown and Lowe, 2007). We find that outputs from 4DiM have higher fidelity than this baseline, and they avoid artifacts like those arising from incorrect exposure adjustment.
@@
We next consider the task of rendering an existing space-time trajectory from novel viewpoints (i.e., video to video translation). We condition on images from a single camera at 8 consecutive frames from a Street View sequence and render a different view yielding 3D-consistent hallucinated regions in the generated images. Results are shown in Fig. 6.
@@ 6 Discussion and future work
To the best of our knowledge, 4DiM is the first model capable of generating multiple, approximately consistent views over simultaneous camera and time control from as few as a single input image. 4DiM achieves state-of-the-art pose alignment and much better generalization compared to prior work on 3D NVS. This new class of models opens the door to myriad downstream applications, some of which we demonstrate, e.g., changing camera pose in videos and seamlessly stitching panoramas. While still not perfect, 4DiM also achieves previously unseen quality in extremely challenging camera trajectories such as rotation from a single image in the wild, indoor or outdoor. We expect 4DiM and follow-up works to have a major impact on methods focused on 3D model extraction, which rely heavily on pose-conditional diffusion models.
@@
Like all generative image and video models, it is important to develop and release models with care. 4DiM is trained largely on data of scenes without people (or anonymized where present), and does not condition on text prompts, which mitigates many of the safety issues that might otherwise arise.
@@
Appendix A Calibrated RealEstate10K
To calibrate RealEstate10K, we predict a single length scale per scene with the help of a recent zero-shot model for monocular depth [Saxena et al., 2023] that predicts metric depth from RGB and FOV. For each image we compute metric depth using the FOV provided by COLMAP. The length-scale is then obtained by regressing the 3D points obtained by COLMAP to metric depth at image locations to which the COLMAP points project. An L1 loss provides robustness to outliers in the depth map estimated by the metric depth models. The mean and variance of the per-frame scales yields a per-sequence estimator. We use the variance as a simple measure of confidence to identify scenes for which the scale estimate may not be reliable, discarding 30% of scenes with the highest variance. Table 6 shows the importance of filtering out scenes with less reliable calibration. The calibrated dataset will be made public to facilitate quantitative comparisons.

### v2

@@ 5 Experiments
| FID () | FDD () | TSEDt=2 () | SfMDpos() | SfMDrot() | LPIPS () | PSNR () | SSIM () | |
| cRE10k test | ||||||||
| PNVS | 51.41 | 1007. | 0.9961 (1.000) | 0.9773 (1.024) | 0.3068 (0.3638) | 0.3899 | 16.07 | 0.3878 |
| MotionCtrl | 43.07 | 370.6 | 0.4193 (1.000) | 1.027 (1.032) | 0.2549 (0.3634) | 0.5003 | 12.74 | 0.2668 |
| ZeroNVS | 37.35 | 522.3 | 0.6638 (1.000) | 1.049 (1.028) | 0.3323 (0.3579) | 0.3792 | 15.30 | 0.3617 |
| 4DiM-R | 31.23 | 306.3 | 0.9974 (1.000) | 1.023 (1.075) | 0.3029 (0.3413) | 0.2630 | 18.09 | 0.5309 |
| 4DiM | 31.96 | 314.9 | 0.9935 (1.000) | 1.008 (1.034) | 0.3326 (0.3488) | 0.3016 | 17.08 | 0.4628 |
| cLLFF zero-shot | ||||||||
| PNVS | 185.4 | 2197. | 0.7235 (0.9962) | 0.9346 (0.8853) | 0.2213 (0.1857) | 0.5969 | 12.04 | 0.1311 |
| MotionCtrl | 106.0 | 531.5 | 0.1515 (0.9962) | 0.9148 (0.9415) | 0.1366 (0.2216) | 0.7016 | 9.722 | 0.06756 |
| ZeroNVS | 75.95 | 860.2 | 0.2205 (0.9962) | 0.9315 (0.8980) | 0.2015 (0.1947) | 0.5008 | 12.58 | 0.1578 |
| 4DiM-R | 63.48 | 353.2 | 0.9659 (0.9962) | 0.9265 (0.8951) | 0.2011 (0.1934) | 0.5403 | 11.55 | 0.1444 |
| 4DiM | 63.78 | 356.8 | 0.9167 (0.9962) | 0.9131 (0.8960) | 0.1838 (0.1943) | 0.5415 | 11.58 | 0.1408 |
| FID () | FDD () | TSEDt=2 () | SfMDpos() | SfMDrot() | LPIPS () | PSNR () | SSIM () | |
| RE10k test | ||||||||
| 4DiM-R (RE10k) | 32.40 | 335.9 | 1.000 (1.000) | 1.021 (1.022) | 0.3508 (0.3737) | 0.3002 | 16.64 | 0.4704 |
| 4DiM-R (cRE10k) | 31.23 | 306.3 | 0.9974 (1.000) | 1.023 (1.075) | 0.3029 (0.3413) | 0.2630 | 18.09 | 0.5309 |
| LLFF zero-shot | ||||||||
| 4DiM-R (RE10k) | 71.07 | 521.2 | 0.7727 (0.9962) | 0.9003 (0.8876) | 0.2108 (0.1915) | 0.4568 | 12.87 | 0.2016 |
| 4DiM-R (cRE10k) | 63.48 | 353.2 | 0.9659 (0.9962) | 0.9265 (0.8951) | 0.2011 (0.1934) | 0.5403 | 11.55 | 0.1444 |
| ScanNet++ zero-shot | ||||||||
| 4DiM-R (RE10k) | 23.68 | 248.0 | 0.9512 (0.9815) | 1.885 (1.858) | 1.534 (1.520) | 0.1938 | 20.74 | 0.6716 |
| 4DiM-R (cRE10k) | 22.89 | 243.2 | 0.9685 (0.9815) | 1.851 (1.917) | 1.541 (1.550) | 0.1809 | 21.25 | 0.6952 |
| FID () | FDD () | TSEDt=2 () | SfMDpos() | SfMDrot() | LPIPS () | PSNR () | SSIM () | |
| cRE10k test | ||||||||
| 4DiM (no video) | 32.45 | 326.5 | 0.9974 (1.000) | 1.023 (1.029) | 0.3356 (0.3692) | 0.2713 | 17.80 | 0.5209 |
| 4DiM | 31.96 | 314.9 | 0.9935 (1.000) | 1.008 (1.034) | 0.3326 (0.3488) | 0.3016 | 17.08 | 0.4628 |
| ScanNet++ test | ||||||||
| 4DiM (no video) | 24.61 | 250.7 | 0.9615 (0.9815) | 1.895 (1.764) | 1.446 (1.431) | 0.1824 | 21.34 | 0.6969 |
| 4DiM | 23.48 | 223.7 | n/a (0.9815) | 1.706 (1.889) | 1.550 (1.503) | 0.1965 | 20.67 | 0.6434 |
| cLLFF zero-shot | ||||||||
| 4DiM (no video) | 64.34 | 401.4 | 0.9886 (0.9962) | 0.9157 (0.9067) | 0.1872 (0.1965) | 0.5437 | 11.45 | 0.1430 |
| 4DiM | 63.78 | 356.8 | 0.9167 (0.9962) | 0.9131 (0.8960) | 0.1838 (0.1943) | 0.5415 | 11.58 | 0.1408 |
@@
We consider both in-distribution and OOD evaluations for 3D NVS in ablations and comparisons to prior work. We use the RealEstate10K dataset (Zhou et al., 2018) as a common in-distribution evaluation. We use 1% of the dataset as our validation split and compute metrics on all baselines ourselves for this split, noting they might instead be advantaged as our test data may exist in their training data (for PNVS, all our test data is in fact part of their training dataset). For the OOD case, we use the LLFF dataset (Mildenhall et al., 2019). We present our main results on 3D novel view synthesis conditioned on a single image in Tab. 1 and Fig. 3, comparing the capabilities of our 4DiM cascade against PhotoConsistentNVS***For PNVS, we follow Yu et al. (2023a) and use a Markovian sliding window for sampling, as they find it is the stronger than stochastic conditioning (Watson et al., 2022). (Yu et al., 2023a) (PNVS), MotionCtrl (Wang et al., 2023b),
@@
and the diffusion model prepared by ZeroNVS (Sargent et al., 2023). These are the strongest 3D NVS diffusion models for natural scenes with code and checkpoints available. For LLFF, we load the input trajectory of views in order, subsample frames evenly, and then generate 7 views given a single input image. RealEstate10K trajectories are much longer, so following PNVS, we subsample with a stride of 10. We use the 7-view NVS task conditioned on a single image, in order to avoid weakening the baselines: MotionCtrl can only predict up to 14 frames, and PNVS performance degrades with the length of the sequence as it is an image-to-image model. We trained versions of 4DiM that process 8-frames to this end (as opposed to using the main 32-frame 4DiM model and subsampling the output) for a more apples-to-apples comparison. Quantitative metrics are computed on 128 scenes from our RealEstate10K test split, and on all 44 scenes in LLFF.
@@
We find that 4DiM achieves superior results in image quality (FID, FDD, FVD) by a wide margin compared to all the baselines, and generally outperforms them in reconstruction-based metrics (PSNR, SSIM, LPIPS) with the exception of ZeroNVS on the LLFF dataset. As much prior work has shown (Saharia et al., 2022c; Watson et al., 2022), reconstruction-based metrics penalize generated samples that may be plausible and tend to favor blurry images (for instance, diffusion architectures trained purely as regression models tend to outperform their generative counterparts in these metrics). Indeed, ZeroNVS generally produces samples with significant blur (more qualitative comparisons are included in our Supplementary Material G). Qualitatively, we find that MotionCtrl does not suffer similarly from blur, but has trouble aligning with the conditioning target poses, as is evident in Figure 3. We also observe that PNVS exhibits more artifacts, with a substantial drop in quality in the zero-shot setting. Surprisingly, PNVS achieves the best TSED in LLFF but upon closer inspection, we find that the number of SIFT keypoint matches is 3x less than for 4DiM when computing TSED scores. This suggests potential improvements for TSED which we leave for future work, such as a higher threshold for the number of minimum keypoint matches, or requiring more spatial coverage of keypoints to discard examples for which the matches are too localized.
@@
One of the most critical ingredients for 3D NVS, especially in the out-of-distribution setting, is the training data mixture. We ablate this in three ways:
@@
Table 1 shows that adding more diverse 3D data beyond RealEstate10K to the training mixture improves pose alignment (SfM distances) on LLFF (zero-shot). The evaluation suggests a very small loss in fidelity when using the full mixture, which is expected since the mixture is more diverse and includes both indoor and outdoor scenes.
@@
To determine the impact of consistent, metric scale training data, we compare a model trained with on the original RE10k data against one with calibrated metric scale poses. Neither model includes the other 3D data sources in their training data used to train our best 4DiM model. Quantitative results on 3D NVS from a single conditioning image are shown in Tab. 2 (including zero-shot performance on LLFF and ScanNet++), and qualitative results in Fig. 4. We find that models trained on uncalibrated data often overshoot or undershoot when scale is ambiguous, and that models trained on scale-calibrated data solve this problem.
@@
We also show the importance of video as a source of training data. Table 3 shows results comparing 8-frame 4DiM trained on our final data mixture to an otherwise identical model but trained without video. We find that including video data improves fidelity and generalization ability which is expected since available 3D/4D data is not as diverse. The improvement is significantly more apparent qualitatively (see Supplementary Material I for side-by-side samples), where it is clear that without video geometry becomes much worse.
@@ 5.2 Emergence of temporal dynamics
| Input | ||||||
StreetView |
| Stitched using gamma adjustment and homography warp | ||||||
| Generated | ||||||
@@
Early in our experiments, we observed that video from a single image can often yield samples with little motion. While this can be remedied by using more guidance weight on timestamps than on the conditioning images (see our Supplementary Material B), the single-image-conditioned case remains difficult. Nevertheless, with two or more input frames, 4DiM can successfully interpolate and extrapolate video and generate very rich dynamics. We quantitatively illustrate these results through our keypoint distance metric (along with other standard metrics) in Tab. 4. We also include quantitative and qualitative comparisons to prior frame interpolation methods in our Supplementary Material H. We suspect that future work including additional controls (e.g., text) on models like 4DiM, as well as larger scale, will likely play a key role in breaking modes for temporal dynamics, given that text+image-to-video has been successful for large models (Brooks et al., 2024b) and the same large-scale efforts report that at smaller scales dynamics quickly worsen (e.g., see SORA samples with 16x less compute).
@@ 5.3 Multiframe Conditioning
While we (and most baselines) focus on single frame conditioning, using more than one frame unlocks interesting capabilities. Here we show two applications. First, we consider 4DiM on the task of panorama stitching (rendering novel viewpoints), given discrete images that collectively cover a full 360 FOV (see Fig. 5). For reference we stitch the images using a naive homography warp and gamma adjustment (Brown & Lowe, 2007). We find that outputs from 4DiM have higher fidelity than this baseline, and they avoid artifacts like those arising from incorrect exposure adjustment.
@@
We finally consider the task of rendering an existing space-time trajectory from novel viewpoints (i.e., pose-conditional video-to-video translation). In Fig. 6, we condition on images from a single camera at eight consecutive frames from a Street View sequence and render different views yielding 3D-consistent hallucinated regions in the generated images. In our website (https://4d-diffusion.github.io), we include many more samples on video-to-video translation to demonstrate this works well in non-street scenes. We show that 4DiM can be used to stabilize the camera in an input video, and also move it along a novel trajectory even when the input video may feature a different camera motion.
@@ 6 Discussion and future work
To the best of our knowledge, 4DiM is the first model capable of generating multiple, approximately consistent views over simultaneous camera and time control from as few as a single input image. 4DiM achieves state-of-the-art pose alignment and much better generalization compared to prior work on 3D NVS. This new class of models opens the door to myriad downstream applications, some of which we demonstrate, e.g., changing camera pose in videos and seamlessly stitching panoramas. While still not perfect, 4DiM also achieves previously unseen quality in extremely challenging camera trajectories such as rotation from a single image in the wild, indoor or outdoor.
@@
Like all generative image and video models, it is important to develop and deploy models responsibly and with care. 4DiM is trained largely on data of scenes without people (or anonymized where present), and does not condition on text prompts, which mitigates many of the safety issues that might otherwise arise.
@@
This supplementary material expands upon the main paper, ‘Diffusion Models for 4D Novel View Synthesis’, by providing in-depth details and analyses. We first delve into the scale calibration process for the RealEstate10K dataset and the scene filtering methodology (Section A). We also discuss the effect of multi-guidance (Section B). Next, we provide a comprehensive overview of the 4DiM architecture, including its UViT backbone, transformer blocks, and camera ray conditioning (Section C). We then detail the training methodology and compute cost (Section D) and elaborate on the evaluation metrics used (Section E). Section F offers details of the implementation of baseline models from previous work, and Section G presents a qualitative comparison of 360∘ scene generation between 4DiM and ZeroNVS. Finally, we showcase the capabilities of 4DiM with an extensive collection of generated samples (Section J). For a more complete collection of samples, please also visit https://4d-diffusion.github.io.
Appendix A More details on calibrating and filtering RealEstate10K
To calibrate RealEstate10K, we predict a single length scale per scene with the help of a recent zero-shot model for monocular depth (Saxena et al., 2023) that predicts metric depth from RGB and FOV. For each image we compute metric depth using the FOV provided by COLMAP. The length-scale is then obtained by regressing the 3D points obtained by COLMAP to metric depth at image locations to which the COLMAP points project. An L1 loss provides robustness to outliers in the depth map estimated by the metric depth models. The mean and variance of the per-frame scales yields a per-sequence estimator. We use the variance as a simple measure of confidence to identify scenes for which the scale estimate may not be reliable, discarding 30% of scenes with the highest variance. Table 6 shows the importance of filtering out scenes with less reliable calibration. The calibrated dataset will be made public to facilitate quantitative comparisons.